OpenStudy (anonymous):

Solve the equation -6p-21=3p-12 A)1 B)3 C)-3 D)-1

OpenStudy (anonymous):

-6p-21=3p-12 -6p=3p+9 -9p=9 p=-1 So d is the answer

OpenStudy (anonymous):

Thank you ^_^

OpenStudy (anonymous):

Your welcome, if it helped plz give a medal thanks :)